# There may be a unique $Q$-point

### v1

@@
Abstract. We show that in the model obtained by iteratively pseudo-intersecting a Ramsey ultrafilter via a length-$\omega_{2}$ countable support iteration of restricted Mathias forcing over a ground model satisfying CH, there is a unique $Q$-point up to isomorphism.
@@ 2 Result
It is consistent that there exists a unique $Q$-point, and this $Q$-point is a Ramsey ultrafilter.
@@
We may assume without loss of generality that $p^{\prime}\Vdash_{\mathbb{M}_{\mathcal{U}}}\forall n\in\omega:\undertilde{c}(n)\subseteq(\underaccent{\dot}{\eta}(n-1),\underaccent{\dot}{\eta}(n)]$. Let $\undertilde{C}$ be a $\mathbb{M}_{\mathcal{U}}$-name for an element of $[\omega]^{\omega}$ such that $p^{\prime}\Vdash_{\mathbb{M}_{\mathcal{U}}}\undertilde{C}=\bigcup\text{range}(\undertilde{c})$. Hence we have
@@
Note that the conclusion of Lemmas 2.4 and 2.5 also holds for each $\langle s^{\prime},x^{\prime}\rangle\geq_{\mathbb{M}_{\mathcal{U}}}\langle s,x_{2}\rangle$. Hence we simply trim $x_{2}$ such that the enumeration of $s\cup x_{2}$ dominates $2^{j+1}$ above $|s|$ and replace $s$ with $s\cup\{\min x_{2}\}$ and $x_{2}$ with $x_{2}\setminus\{\min x_{2}\}$.   $\dashv$
@@
Since $l_{0}$ is the $(j-1)$’th element of $s\cup t^{\prime}$, we have $l_{0}>2^{j}$ by Fact 2.6.333Note that the additional requirement in Fact 2.6 that $\max s$ is already larger than $2^{|s|}$ is needed here, since $l_{0}$ could be $\max s$. Hence, since $l_{0}\in N_{i^{*}}$, it follows that $\gamma\in N_{i^{*}}$. Thus, all the parameters in the above formula lie in $N_{i^{*}}$, which implies that there exists $\langle m^{\dagger},l^{\dagger},D^{\dagger}\rangle\in N_{i^{*}+1}$ satisfying the formula.

### v2

# A unique $Q$-point and infinitely many near-coherence classes of ultrafilters

@@
Abstract. We show that in the model obtained by iteratively pseudo-intersecting a Ramsey ultrafilter via a length-$\omega_{2}$ countable support iteration of restricted Mathias forcing over a ground model satisfying CH, there is a unique $Q$-point up to isomorphism. In particular, it is consistent that there is only one $Q$-point while there are $2^{\mathfrak{c}}$-many near-coherence classes of ultrafilters.
@@ 1 Introduction
While the consistency of the non-existence of $Q$-points is a well-established fact with a variety of witnesses apart from the Mathias model333such as the Laver and Miller models (see [7] and [11], respectively)., the construction of models containing only ‘few’ $Q$-points seems to have received less attention. However, such models do arise naturally as models containing only few near-coherence classes of ultrafilters [^4] Two ultrafilters $\mathcal{U}_{1}$ and $\mathcal{U}_{2}$ are nearly-coherent iff there is some finite-to-one $f\in{{}^{\omega}}{\omega}$ such that $f(\mathcal{U}_{1})=f(\mathcal{U}_{2})$, where $f(\mathcal{U}_{i}):=\{X\subseteq\omega:f^{-1}[X]\in\mathcal{U}_{i}\}$. Note that two $Q$-points are nearly-coherent iff they are isomorphic.: Indeed, Mildenberger [10] has constructed models with exactly two and exactly three near-coherence classes, and it is easy to see that these contain exactly one and exactly two $Q$-points, respectively: In her model with exactly two near-coherence classes, one class contains a Ramsey ultrafilter, while the other class contains an ultrafilter that is $\omega_{1}$-generated. Hence, this latter class cannot contain a $Q$-point, since her models satisfy $\mathfrak{d}=\mathfrak{c}=\omega_{2}$ and such a $Q$-point would thus have to be ${<}\mathfrak{d}$-generated, which is impossible. Analogously, in Mildenberger’s model with exactly three near-coherence classes, two classes are represented by Ramsey ultrafilters, while the third contains an $\omega_{1}$-generated ultrafilter – giving exactly two $Q$-points in total.
The construction of models with exactly $n$ near-coherence classes of ultrafilters for various finite $n\geq 4$ would similarly yield the consistency of exactly $m$ $Q$-points for some $m<n$ [^5] The inequality is strict since such a model must satisfy $\mathfrak{u}<\mathfrak{d}$, a result due to Banakh and Blass [9]. Hence, one of the $n$ near-coherence classes contains a ${<}\mathfrak{d}$-generated ultrafilter and thus no $Q$-point.
The model considered in this paper is of a different nature, however: It contains only one $Q$-point while its number of near-coherence classes is $2^{\mathfrak{c}}$, i.e., the model’s lack of $Q$-points is not the consequence of a lack of near-coherence classes. This follows from the fact that dominating reals are added at each of the $\omega_{2}$-many stages of the iteration, which gives $\mathfrak{b}=\mathfrak{d}=\mathfrak{c}=\omega_{2}$ in the final extension. Since $\mathfrak{b}\leq\mathfrak{u}$ (see Solomon [8]), we have $\mathfrak{u}=\mathfrak{d}=\omega_{2}$ in our model, and hence there are $2^{\mathfrak{c}}$-many near-coherence classes of ultrafilters by Banakh and Blass [9].
@@ 2 Result
It is consistent that there is a unique $Q$-point while there are $2^{\mathfrak{c}}$-many near-coherence classes of ultrafilters.
@@
We may assume without loss of generality that $p^{\prime}\Vdash_{\mathbb{M}_{\mathcal{U}}}\forall n\in\omega:\undertilde{c}(n)\subseteq(\underaccent{\dot}{\eta}(n-1),\underaccent{\dot}{\eta}(n)]$. Let $\undertilde{C}$ be a $\mathbb{M}_{\mathcal{U}}$-name for an element of $[\omega]^{\omega}$ such that $p^{\prime}\Vdash_{\mathbb{M}_{\mathcal{U}}}\undertilde{C}=\bigcup\text{range}(\undertilde{c})$. Hence, we have
@@
Note that the conclusion of Lemmas 2.4 and 2.5 also holds for each $\langle s^{\prime},x^{\prime}\rangle\geq_{\mathbb{M}_{\mathcal{U}}}\langle s,x_{2}\rangle$. Hence, we simply trim $x_{2}$ such that the enumeration of $s\cup x_{2}$ dominates $2^{j+1}$ above $|s|$ and replace $s$ with $s\cup\{\min x_{2}\}$ and $x_{2}$ with $x_{2}\setminus\{\min x_{2}\}$.   $\dashv$
@@
Since $l_{0}$ is the $(j-1)$’th element of $s\cup t^{\prime}$, we have $l_{0}>2^{j}$ by Fact 2.6.666Note that the additional requirement in Fact 2.6 that $\max s$ is already larger than $2^{|s|}$ is needed here, since $l_{0}$ could be $\max s$. Hence, since $l_{0}\in N_{i^{*}}$, it follows that $\gamma\in N_{i^{*}}$. Thus, all the parameters in the above formula lie in $N_{i^{*}}$, which implies that there exists $\langle m^{\dagger},l^{\dagger},D^{\dagger}\rangle\in N_{i^{*}+1}$ satisfying the formula.